OpenStudy (anonymous):

let f(x)=-2x +4 and g(x) = -6x - 7 find f(x) - g(x) ... HELP PLEASE

OpenStudy (amonoconnor):

Okay, this is what it'll look like: f(x) - g(x) (2x + 4) - (-6x - 7) 2x + 4 + 6x + 7 = 8x + 11

OpenStudy (amonoconnor):

Does that make sense? :)

OpenStudy (anonymous):

is that the answer?

OpenStudy (amonoconnor):

Yep! Unless the gave you a number to plug in for x, that's all you can do. :)

OpenStudy (anonymous):

this is the question i didnt know how to get the circle thing in there i hope you aren't confused like i am

OpenStudy (anonymous):

does that change anything?? or is that still the answer

OpenStudy (amonoconnor):

Do you have options of answers to choose from?

OpenStudy (anonymous):

Nope :( i have to write out how to explain it

OpenStudy (amonoconnor):

Is this a second question, or the one you posted... The equations don't look the same (?)

OpenStudy (anonymous):

oh shoot i have two questions that are alike my bad hold on :p

OpenStudy (amonoconnor):

Okay. :)

OpenStudy (amonoconnor):

Yep, this is like my original post. (-2x + 4) - (-6x - 7) -2x + 4 + 6x + 7 = 4x + 11

OpenStudy (amonoconnor):

Do you want the other one too?

OpenStudy (anonymous):

yes please <3 i appreciate this so much honestly!

OpenStudy (amonoconnor):

Oh, and I guess an explanation is... "Using properties of addition and subtraction, I combined like terms."

OpenStudy (amonoconnor):

Okay. I'm on it :)

OpenStudy (anonymous):

thank you :D <3

OpenStudy (amonoconnor):

So... (3x - 7)(-2x - 6) (3(4) - 7)(-2(4) - 6) (12 - 7)(-8 - 6) (5)(-14) = -70

OpenStudy (amonoconnor):

Yep, that's it. Checked it on paper :)

OpenStudy (amonoconnor):

Thanks for the medal ;)

OpenStudy (amonoconnor):

Any more?

OpenStudy (anonymous):

No, but thank you so much you saved my butt :O

OpenStudy (amonoconnor):

You bet!

OpenStudy (anonymous):

<3 :D